A man wants to coat the ceiling, walls, and floor of a rectangular room with a coating that cost $39.95 per gallon. The dimensions of the room are 7m by 9m by 4.5m. If one gallon covers 110m/2, how much will the man spend to cover the room?
Answer by Alan3354(69443)

A man wants to coat the ceiling, walls, and floor of a rectangular room with a coating that cost $39.95 per gallon. The dimensions of the room are 7m by 9m by 4.5m. If one gallon covers 110m/2, how much will the man spend to cover the room?
----------------
Find the area of the room. It has 6 sides that are 3 pairs.
A = 2*(7*9) + 2*(7*4.5) + 2*(9*4.5)
A = 126 + 63 + 81
A = 270 sq meters
1 gallon covers 100 sq m, so he'll need 270/110 gallons,
27/11 gallons.
(27/11)*39.95 = $98.06